OpenStudy (anonymous):

Given the triangle below, what is cot∡B? Right triangle ABC with AC measuring 4 and BC measuring 7 square root of 33 over 4 7 times square root of 33 over 33 square root of 33 over 7 4 times square root of 33 over 33

OpenStudy (anonymous):

this one?

OpenStudy (anonymous):

yess Imma Post the Picture (:

OpenStudy (anonymous):

k

OpenStudy (anonymous):

No /:

OpenStudy (anonymous):

do you know how to use the pythagorean theorem to find the missing side of the triangle?

OpenStudy (anonymous):

Lol Nope But my older brother thinks its square root of 33 over 7 is he right ?

OpenStudy (anonymous):

no.... it's not \(\large \frac{\sqrt{33}}{7} \) do you know where the \(\large \sqrt{33} \) came from though?

OpenStudy (anonymous):

Sadly No /.\

OpenStudy (anonymous):

Nvm My Sister Explained It Thanks Though (:

OpenStudy (anonymous):

ok... the sqrt(33) came from the pythagorean theorem: \(\large a^2+4^2=7^2 \) solving that you'll get \(\large a=\sqrt{33} \) but now the trigonometry part... cotangent is adjacent over opposite... ok.... nvmd..